OpenStudy (anonymous):

Mindy is taller than Olga, but shorter than Thomas. Thomas is taller than Sven, but shorter than Parth. Sven is taller than Olga, but shorter than Mindy. Who is the tallest?

TheSmartOne (thesmartone):

Mindy is older than Olga so Olga is definitely not the tallest

TheSmartOne (thesmartone):

But hen Thomas is taller than Mindy so Mindy is not the tallest.

TheSmartOne (thesmartone):

hen should be then

OpenStudy (anonymous):

Shouldn't it be Parth?

TheSmartOne (thesmartone):

It says Thomas is taller than Sven so Sven is out

OpenStudy (jdoe0001):

\(\large \begin{array}{llll} mindy>olga\\ mindy<thomas \end{array}\qquad \begin{array}{llll} thomas>sven\\ thomas < parth \end{array}\qquad \begin{array}{llll} sven > olga\\ sven < mindy \end{array}\)

TheSmartOne (thesmartone):

But then it says Thomas is shorter than Parth so it should be Parth who is tallest

OpenStudy (jdoe0001):

yeap \(\large { \begin{array}{llll} mindy>olga\\ mindy<thomas \end{array}\qquad \begin{array}{llll} thomas>sven\\ thomas < parth \end{array}\qquad \begin{array}{llll} sven > olga\\ sven < mindy \end{array} \\ \quad \\ olga < sven < mindy < thomas < parth }\)

TheSmartOne (thesmartone):

Everyone agrees the Parth is the tallest.. @amandawells983